### v3

# Analyzing the Performance Portability of SYCL across CPUs, GPUs, and Hybrid Systems with SW Sequence Alignment

Abstract
The high-performance computing (HPC) landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments. This comprehensive study extends our previous research on SYCL’s performance portability by evaluating its effectiveness across a broader spectrum of computing architectures, including CPUs, GPUs, and hybrid CPU-GPU configurations from NVIDIA, Intel, and AMD. Our analysis covers single-GPU, multi-GPU, single-CPU, and CPU-GPU hybrid setups, using two common, bioinformatic applications as a case study. The results demonstrate SYCL’s versatility across different architectures, maintaining comparable performance to CUDA on NVIDIA GPUs while achieving similar architectural efficiency rates on AMD and Intel GPUs in the majority of cases tested. SYCL also demonstrated remarkable versatility and effectiveness across CPUs from various manufacturers, including the latest hybrid architectures from Intel. Although SYCL showed excellent functional portability in hybrid CPU-GPU configurations, performance varied significantly based on specific hardware combinations. Some performance limitations were identified in multi-GPU and CPU-GPU configurations, primarily attributed to workload distribution strategies rather than SYCL-specific constraints. These findings position SYCL as a promising unified programming model for heterogeneous computing environments, particularly for bioinformatic applications.
keywords:
SYCL, GPU, CPU-GPU, oneAPI, CUDA, Performance portability, Heterogeneous computing
journal: Future Generation Computer Systems
[…]
[L] © $<$2025$>$. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/
The published journal article is available at: https://doi.org/10.1016/j.future.2025.107838
[…]
[fn1]organization=III-LIDI, Facultad de Informática, addressline=50 y 120, city=La Plata, postcode=1900, state=Buenos Aires, country=Argentina
[…]
[fn2]organization=Comisión de Investigaciones Científicas (CIC), city=La Plata, state=Buenos Aires, country=Argentina
[…]
[fn3]organization=Dpto. Arquitectura de Computadores y Automática, Universidad Complutense de Madrid, city=Madrid, postcode=28040, country=España
1 Introduction
The parallel processing landscape is undergoing rapid transformation, with an increasing emphasis on energy-efficient and heterogeneous computing environments [navaux23]. GPUs are now the dominant accelerator in this field, offering incomparable performance for parallel workloads. Still, the roles of CPUs and the potential of hybrid CPU-GPU systems are also actively explored in computational research. On the one hand, GPU-based Systems on Chip (SoC) are becoming more popular every day. On the other hand, hybridization now extends to CPU architectures, with the emergence of heterogeneous designs like Intel’s Raptor Lake and Meteor Lake. Depending on the workload, these processors incorporate up to three different types of cores on the same die to optimize performance and energy efficiency. This architectural heterogeneity at the system and processor levels reflects the larger market dynamics and the ongoing evolution of computing architectures.
As of 2024, NVIDIA maintains its dominance in the GPU market, particularly in the discrete GPU (dGPU) segment, where it possesses an 88% market share. AMD has demonstrated significant progress, capturing 12% of the market, while Intel experienced a decline, with its market share dropping to less than 1% [shilov24]. In the integrated graphics (iGPU) sector, Intel retains a substantial advantage with a 68% market share, followed by NVIDIA (18%) and AMD (14%). This market structure highlights the importance of considering diverse architectures in GPU-based computing research [norem24].
Concerning parallel programming, the proliferation of diverse computing architectures (including CPUs, GPUs, and hybrid CPU-GPU systems), has created a pressing need for unified programming models that can effectively harness their collective potential. In this context, SYCL has emerged as a promising solution that offers a cross-platform abstraction layer that facilitates code portability across various hardware architectures. By enabling developers to write high-performance code that can be seamlessly executed on a wide range of devices, SYCL plays a crucial role in addressing the challenge of performance portability in heterogeneous computing environments [SYCL].
This article evolves from our systematic exploration of SYCL’s capabilities in bioinformatics applications across heterogeneous computing platforms. In our initial work [Costanzo2022IWBBIO], we focused on migrating a CUDA-based, protein database search software (SW#db) to SYCL using Intel’s oneAPI ecosystem. This study demonstrated the tool’s effectiveness for code migration and it defined the basic cross-GPU-architecture portability of the migrated code, achieving comparable or slightly better efficiency rates compared to the native CUDA implementation. Building upon these findings, in [costanzo23] we expanded the study to evaluate SYCL’s performance portability across different GPU architectures, including single and multi-GPU configurations from different vendors. This research revealed that while CUDA and SYCL versions achieved similar performance on NVIDIA devices, SYCL reached higher architectural efficiency rates in most test cases from the other GPU vendors. Finally, in [supercompcost23] we completed the SYCL-migration process of SW# biological suite to support pairwise alignment and different alignment algorithms. Moreover, its performance in a wide variety of biological test scenarios was analyzed and cross-SYCL-implementation portability was verified on several GPUs and some CPUs from a functional perspective.
This work extends our performance portability study to include pairwise sequence alignment and (pure and hybrid) CPUs and CPU-GPU systems. The key contributions of this paper are:
  • 1.
    An extension of our previously adapted performance model from [lan2017swhybrid], to encompass pure and hybrid CPU architectures and also combined CPU-GPU systems.
  • 2.
    An extension of our previous work on the functional and performance portability of SW# applications across different GPU architectures to now include pairwise sequence alignment and Intel and AMD CPUs, as well as hybrid CPU-GPU configurations. On the one side, by investigating the performance characteristics of SYCL on these diverse platforms, this study aims to provide valuable insights into its potential as a unified programming model for the increasingly complex landscape of modern computing. On the other side, by expanding the study to include CPUs and hybrid systems, this research seeks to offer a more comprehensive understanding of how SYCL-based applications perform across the full spectrum of available computational resources.
  • 3.
    An analysis of the trade-offs and synergies between CPU and GPU computations in hybrid configurations for two common bioinformatic workloads.
The remainder of this paper is structured as follows: Section 2 presents the background; Section 3 details the case study applications and the expanded performance model; Section 4 presents our findings on functional and performance portability; Section LABEL:sec:relworks discusses related work; and Section LABEL:sec:conc presents the conclusions and future research directions.
2 Background
2.1 CPUs, GPUs, and Programming Models for Heterogeneous Computing
The evolution of high-performance computing (HPC) has been significantly influenced by the emergence of general-purpose GPU programming. In 2007, NVIDIA introduced CUDA [CUDA_handson] alongside the Tesla GPU architecture, enabling non-graphics applications to harness GPU computational power. Although CUDA quickly became a prominent low-level programming model for GPU computing, its proprietary nature and exclusive compatibility with NVIDIA hardware led to the development of more versatile alternatives, such as OpenCL [OpenCL]. OpenCL offers a level of abstraction comparable to that of CUDA while supporting multiple devices and vendors.
The field has also expanded to include multicore CPUs and hybrid GPU-CPU systems, accompanied by the development of high-level programming initiatives such as OpenMP [OpenMP], OpenACC [OpenACC, OpenACC_book], and SYCL [SYCL]. Although CUDA is limited to NVIDIA hardware, and OpenCL provides broad coverage but is challenging to learn and costly to implement, SYCL has emerged as a promising solution for heterogeneous computing.
SYCL [SYCL] enables developers to write code for various processors, including CPUs, GPUs, and hybrid systems, using standard ISO C++. It incorporates host and device code in a single source file and supports multiple acceleration APIs, such as OpenCL, facilitating seamless integration with lower-level code across different hardware architectures.
Several SYCL implementations are now available, with Intel’s oneAPI [DPCPP] standing out as a mature developer suite. OneAPI eliminates the need for separate code bases for host and device, which is required in OpenCL, and supports multiple programming languages and tools for different architectures. Data Parallel C++ (DPC++), oneAPI’s core language [DPCPP], integrates the SYCL and OpenCL standards without additional extensions, facilitating the development of applications that can efficiently utilize the full spectrum of available computational resources.
2.2 Smith-Waterman Sequence Alignment
A key process in bioinformatics and computational biology is sequence alignment, which aims to identify regions of similarity between sequences to uncover their structural, functional, and evolutionary relationships [PSAreview]. The Smith-Waterman (SW) algorithm is commonly employed to determine the optimal local alignment between two sequences [Smith1981]. This technique relies on dynamic programming and is highly sensitive, as it examines all possible alignments between the sequences.
Given two sequences $Q$ and $D$ of length $|Q|=m$ and $|D|=n$, the recurrence relations for the SW algorithm with the modification of Gotoh [gotoh81] are defined as follows:
$H_{i,j}=max\begin{cases}0\\ H_{i-1,j-1}+SM(Q[i],D[j])\\ E_{i,j}\\ F_{i,j}\end{cases}$ (1)
$E_{i,j}=max\begin{cases}H_{i,j-1}-G_{o}\\ E_{i,j-1}-G_{e}\end{cases}$ (2)
$F_{i,j}=max\begin{cases}H_{i-1,j}-G_{o}\\ F_{i-1,j}-G_{e}\end{cases}$ (3)
The similarity score $H_{i,j}$ is calculated to identify a common subsequence; $H_{i,j}$ contains the score to align the prefixes $Q[1..i]$ and $D[1..j]$. Moreover, $E_{i,j}$ and $F_{i,j}$ correspond to the scores of prefix $Q[1..i]$ and $D[1..j]$ aligned to a gap, respectively. SM denotes the scoring matrix and defines the match/mismatch scores between the residues. Lastly, $G_{o}$ and $G_{e}$ refer to the gap open and gap extension penalties, respectively.
Firstly, $H$, $E$ and $F$ must be initialized with 0 when $i=0$ or $j=0$. Then, the recurrences should be calculated with $1\leq i\leq m$ and $1\leq j\leq n$. The highest value in the matrix $H$ ($S$) corresponds to the optimal local alignment score between $Q[1..i]$ and $D[1..j]$. If required, the optimal local alignment is finally obtained by following a traceback procedure whose starting point is $S$. From a computational perspective, it is important to highlight the dependencies of any $H$ element. Any cell can be calculated only after the values of the upper, left, and upper-left neighbors are known; imposing restrictions on how H can be processed.
SW is usually employed to compute: (a) pairwise alignments of two very long DNA sequences, and (b) database similarity search (one-to-many), involving shorter protein sequences. As SW runs in quadratic time to compute optimal alignment, widely adopted parallelization approaches are intra-sequence (a single matrix is calculated and all Processing Elements (PEs) work collaboratively) and inter-sequence (simultaneously calculating multiple matrices without communication between the PEs). Fig. 1 illustrates both approaches.
Figure 1: Parallelization approaches in similarity matrix computations (adapted from [swipe11]). Each color indicates the cells that can be computed together in a SIMD manner.
2.3 Performance portability
According to Penycook et al. [performance_portability_paper], performance portability refers to ”A measurement of an application’s performance efficiency for a given problem that can be executed correctly on all platforms in a given set”. These authors define two different performance efficiency metrics: architectural efficiency and application efficiency. The former denotes the capacity of an application to effectively utilize hardware resources, measured as a proportion of the theoretical peak performance. The latter signifies the application’s ability to select the most suitable implementation for each platform, representing a fraction of the highest observed performance achieved.
[…]
where $e_{i}(\alpha,p)$ corresponds to the performance efficiency of case-study application $\alpha$ solving problem p on the platform i.
The performance portability concept emphasizes the capability to write code that can efficiently utilize the available computing resources, such as CPUs, GPUs, or specialized accelerators, while maintaining high performance regardless of the specific hardware configuration. With performance portability, developers can write code once and have it deliver optimal performance on various target platforms. This eliminates the need for extensive manual code optimizations or platform-specific modifications, reducing development time and effort.
3 Case-Study Applications and Performance Model
3.1 Case-Study Applications
Two GPU-accelerated implementations of SW sequence alignment were considered for the performance portability evaluation:
  • 1.
    CUDA: The SW# framework represents a CUDA-implemented solution for biological sequence alignment that prioritizes memory efficiency. Operating as both a standalone program and an integrated library, it enables protein and DNA sequence comparisons through pairwise alignment and database searches. The framework provides customizable alignment parameters, including Smith-Waterman algorithm options, gap penalty specifications, and scoring matrix selection. Through CPU-GPU coordination, SW# achieves good performance by implementing a dynamic load balancing system that assigns tasks based on sequence length characteristics. The parallelization architecture primarily leverages GPU capabilities through both inter-task and intra-task approaches. GPU processing follows a dual-kernel strategy: shorter database sequences are processed via inter-task parallelism in a short kernel, while longer sequences utilize intra-task parallelism in a long kernel. For multi-GPU configurations, SW# employs an adaptive execution protocol. When the GPU count exceeds query sequence numbers, all devices synchronously process different database segments for a single query. Alternatively, with more queries than GPUs, each GPU independently processes distinct queries against the complete database [swsharp, swsharpdb].
  • 2.
    SYCL: this code is based on the implementation presented in the paper [supercompcost23], representing a SYCL equivalent. The migration of the SW# suite was performed using dpct (the Data Parallel Compatibility Tool available in the oneAPI suite) and some hand-coding modifications. The resulting implementation is pure SYCL code without device-specific optimizations or native code paths, ensuring that all computations—whether on GPUs or CPUs—are executed through SYCL. This approach maintains complete portability across different architectures while allowing for a fair comparison of SYCL’s performance in various computing environments.
3.2 Performance Model
Peak theoretical hardware performance must be estimated for all selected devices in this study to compute the performance portability metric. This step requires considering both hardware and algorithm features. In our previous work [costanzo23], we have adapted the performance model from [lan2017swhybrid] to the SW# algorithm and several GPUs. For the present investigation, we extend this adapted model to encompass CPUs from various vendors. In fact, the presented equations can serve as a generic framework for estimating the theoretical peak performance of multiple devices, including CPUs, GPUs, and even FPGAs. The computing capability of different devices can be estimated using Eq. 4:
$Capability=Clock\_Rate\times Throughput\times Lanes$ (4)
where Clock_rate denotes the device’s clock frequency, Throughput represents the number of instructions the device can execute per clock cycle, and Lanes indicates the quantity of SIMD vector lanes. Subsequently, it is imperative to enumerate the instructions issued during each cell update of the similarity matrix. In the sequence alignment context, the predominant performance metric is Cell Updates Per Second (CUPS), representing the time for a complete computation of one cell in the similarity matrix (H). The GCUPS (billion CUPS) value is calculated by:
$GCUPS=\frac{Q\times D}{t\times 10^{9}}$ (5)
where $|Q|$ is the total number of residues in the query sequence, $|D|$ is the total number of residues in the database and t is the runtime in seconds [RucciBDAG2016]. Consequently, the theoretical peak GCUPS of any device can be modeled using Equation 6:
$Theo\_peak=\frac{Capability}{Instruction\_count\_one\_cell\_update}$ (6)
Table 1 provides a comprehensive summary of the theoretical peak performance for selected GPUs, calculated using Equation 6. Additionally, the performance metrics for CPUs are elucidated in Table 2. Further detailed analysis and discussion are presented in the subsequent sections of this paper.
3.2.1 SW# core instructions
SW# computes the similarity matrix using 32-bit integers and performs 12 instructions per cell update. Algorithm 1 presents the snippet of cell update in similarity matrix as in Eq. 1, Eq. 2, and Eq. 3. Just adding, subtracting, and maximum instructions are required to perform a single-cell update.
1:$E^{1}=E_{l}-G_{e}$triangleright $E_{l}$: E of its left neighbor
2:$E^{2}=H_{l}-G_{o}$triangleright $H_{l}$: H of its left neighbor
3:$E=max(E^{1},E^{2})$
4:$F^{1}=F_{u}-G_{e}$triangleright $F_{u}$: F of its upper neighbor
5:$F^{2}=H_{u}-G_{o}$triangleright $H_{u}$: H of its upper neighbor
6:$F=max(F^{1},F^{2})$
7:$H=H_{ul}+SM$triangleright $H_{ul}$: H of its upper-left neighbor
8:$H=max(H,E)$
9:$H=max(H,F)$
10:$H=max(H,0)$
11:$A=H$triangleright $A$: an auxiliary variable
12:$S=max(H,S)$triangleright $S$: optimal score
Algorithm 1 Core instructions per cell update in similarity matrix
3.2.2 Architectural features on NVIDIA GPUs
The # Cores in an NVIDIA GPU refers to the number of Streaming Multiprocessors. CUDA does not strictly follow a SIMD execution model but it adopts a similar one denoted as the SIMT model. A warp comprises a group of 32 threads that execute the same instruction stream. According to [lan2017swhybrid], “a warp in SIMT is equivalent to a vector in SIMD, and a thread in SIMT is equivalent to a vector lane in SIMD”. The instruction throughput depends on the CUDA Compute Capability (CC) of each NVIDIA GPU  [^7] https://docs.nvidia.com/cuda/cuda-c-programming-guide/#maximize-instruction-throughput.
3.2.3 Architectural features on AMD GPUs
In the RDNA2.0 architecture, the # Cores parameter corresponds to the number of Compute Units (CUs), which are organized in pairs to form Workgroup Processors (WPs). AMD employs the terms wavefront and work-item as analogs to NVIDIA’s warp and thread, respectively. While RDNA2.0 supports wavefront sizes of both 32 and 64 work items, the former has priority over the latter. Each CU encompasses two SIMD32 vector units, enabling the execution of 64 add/subtract/max instructions per cycle (Int32). Consequently, the instruction throughput per work item is quantified as 2.
3.2.4 Architectural features on Intel GPUs
In the dGPU segment, Intel’s GPU design philosophy diverges significantly from that of NVIDIA and AMD. The fundamental building block of the Intel Xe microarchitecture is the Xe Core, each comprising 16 Xe Vector Engines (XVEs)  [^8] Also referred to as Execution Units (EUs), capable of executing 8 add/subtract/max instructions per cycle (Int32). In the context of the proposed model, Xe Cores and XVEs correspond to # Cores and # Lanes, respectively.
In the iGPU segment, both Gen9 and Gen12 microarchitectures exhibit similar design principles, primarily differentiated by the number of computational resources. These microarchitectures utilize the Subslice as their fundamental block, each containing 8 Execution Units (EUs) that can process 8 add/subtract/max instructions per cycle (Int32). Within the framework of the proposed model, Subslices and EUs are analogous to # Cores and # Lanes, respectively.
3.2.5 Architectural features on CPUs
For Intel and AMD CPUs, # Cores represents the number of processor cores, each capable of independently executing instructions, which is crucial for performance in parallel applications. The # Lanes is defined by the number of Vector Processing Units (VPUs) and their supported vector width/instruction set (such as SSE, AVX, AVX-512), which allows a core to perform multiple operations simultaneously. The instruction throughput depends on the operation that is being performed.
Since intensive use of VPUs leads to increased heat dissipation and power consumption, modern processors incorporate various technologies that reduce the clock frequency to counteract their impact [Rutter]. For this reason, the operating frequency when all cores perform intensive computation is usually the base frequency or a slightly higher one, particularly on the server segment. For this work, we have monitored the SW# execution on each CPU device to verify the information provided.
Finally, a caveat must be made for hybrid CPU architectures (such as Intel’s Alder Lake) [AlderLakes]. As these architectures present up to three types of cores (P-cores, E-cores, and LP E-cores), the maximum theoretical performance is obtained by summing up the partial maximum performances.
Vendor NVIDIA Intel AMD
Model GTX 980 GTX 1080 RTX 2070 V100 RTX 3070 RTX 3090 Arc A770 UHD 630 UHD 770 Xe-LPG 128EU RX 6700 XT Vega 6
Type Discrete Discrete Integrated Discrete Integrated
Architecture
Maxwell
(CC 5.2)
Pascal
(CC 6.1)
Turing
(CC 7.5)
Volta
(CC 7.0)
Ampere
(CC 8.6)
Ampere
(CC 8.6)
Xe Gen 9.5 Gen 12 Xe RDNA 2.0 GCN5
# Cores 16 20 36 80 46 82 32 3 4 8 40 6
# Lanes 32 16 8 16 32
Inst. throu.
4/2 4/2 2 8 2
Clock (MHz) 1216 1733 1620 1380 1725 1695 2400 1200 1650 2250 2581 1100
Peak (GCUPS)
155.6 277.3 311 588.8 423.2 741.2 819.2 19.2 35.2 192 550.6 35.2
The instruction throughput for GTX 980 and GTX 1080 is 4 for add/subtract and 2 for max/min. The core instructions include 5
add/subtract and 6 max. Thus, the equivalent throughput is 3.
The core instruction count for each cell update is 12.
Table 1: GPU specifications and their theoretical peak performance in terms of GCUPS
Vendor Intel AMD
Model
Core
i5-7400
Core
i5-10400F
Xeon
E5-1620 v3
Xeon
E5-2695 v3
Xeon
Gold 6138
Core
i9-9900K
Core
i9-13900K
Core
Ultra 9-185H
Ryzen 3
5300U
Segment Desktop Server Desktop Mobile
Architecture Kaby Lake Comet Lake
Sandy
Bridge-E
Haswell Skylake
Coffee
Lake-R
Raptor
Lake-S
Meteor
Lake-S
Lucienne
# Cores 4 6 4 14 40 8 8/16 6/8/2 4
SIMD set AVX2 AVX2 AVX2 AVX2 AVX-512 AVX2 AVX2 AVX2 AVX2
# Lanes 8 16 8
Inst. throu. 1
Clock (MHz) 3300 4000 3500 1900 1900 4700 5500/4300 5100/3800/2500 3600
Peak (GCUPS) 8.8 16 9.3 35.5 101.3 25.1 75.2 44 9.6
  • a
    The Intel Core i9-13900K CPU features 8 P-cores and 16 E-cores.
  • b
    The Intel Core Ultra 9-185H CPU features 6 P-cores, 8 E-cores, and 2 LP E-cores.
Table 2: CPU specifications and their theoretical peak performance in terms of GCUPS
4 Experimental Results
4.1 Experimental Design
The experiments were carried out on 12 GPUs, including 6 NVIDIA dGPUs, 1 AMD dGPU, 1 AMD iGPU, 3 Intel iGPUs, and 1 Intel dGPU. In addition, 8 Intel CPUs (from different segments) and 1 AMD CPU were included in the study. The specific details of these platforms can be found in Table 1. The oneAPI and CUDA versions used were 2022.1.0 and 11.7, respectively. For both CUDA and SYCL, the optimization flag -O3 was used during compilation. To run SYCL code on NVIDIA and AMD GPUs, several modifications had to be made to the build process, as SYCL is not supported by default on these platforms [^9] https://intel.github.io/llvm-docs/GetStartedGuide.html but Codeplay recently announced free binary plugins [^10] https://codeplay.com/portal/blogs/2022/12/16/bringing-nvidia-and-amd-support-to-oneapi.html to support it. After these modifications, it was possible to run DPC++ code on both NVIDIA and AMD GPUs using the Clang++ compiler (16.0).
To ensure that the CPU executes only SYCL code and avoid the use of SW# “native” CPU code, the flag T=0 was used in all tests. This setting forces all sequence alignments to be thoroughly computed using SYCL, allowing for a fair evaluation of SYCL performance on the CPU.
Performance evaluation was carried out using real biological data for the two application benchmarks:
  • 1.
    Protein database search: searching 20 query protein sequences against the well-known Environmental Non-Redundant database (Env. NR) (2021_04 Release), which contains 995210546 amino acid residues in 4789355 sequences, with a maximum length of 16925. Query sequences were selected from the Swiss-Prot database  [^11] Swiss-Prot:  https://www.uniprot.org/downloads, with lengths ranging from 144 to 5478. The access numbers for these queries are: P02232, P05013, P14942, P07327, P01008, P03435, P42357, P21177, Q38941, P27895, P07756, P04775, P19096, P28167, P0C6B8, P20930, P08519, Q7TMA5, P33450, and Q9UKN1. Besides, SW# was configured with BLOSUM62 as the substitution matrix, and 10/2 as the penalty for insertion/extension.
  • 2.
    Pairwise alignment: computing five different DNA alignments. Table 3 presents the accession numbers and sizes of the DNA sequences used. The score parameters were configured as +1 for match, -3 for mismatch, -5 for gap open, and -2 for gap extension.
Query sequence Database sequence
Accession Size Accession Size
Matrix Size
NC_000898 162K NC_007605 172K 28M
NC_003064.2 543K NC_000914.1 536K 291M
CP000051.1 1M AE002160.2 1M 1G
BA000035.2 3M BX927147.1 3M 9G
NC_005027.1 7M NC_003997.3 5M 35G
Table 3: DNA sequence information used in the tests
Last, each test was executed 20 times and the performance was determined based on the average of these multiple runs (to minimize deviations).
4.2 Single-GPU Performance and Portability Results
A comprehensive analysis of SYCL and CUDA implementations was conducted on various NVIDIA, AMD, and Intel GPUs to evaluate the performance and portability of SYCL across different GPU architectures.
4.2.1 Protein Database Search
Table LABEL:tab:arch-eff provides a detailed comparison of the performance and the architectural efficiency of CUDA and SYCL implementations on each platform for protein database search. On NVIDIA GPUs, CUDA and SYCL demonstrated comparable performance and efficiency values. As anticipated, more powerful GPUs achieved higher GCUPS values. The architectural efficiency values ranged from 37% to 52%. In particular, while the RTX 3090 GPU exhibited the highest GCUPS value, the RTX 2070 GPU proved to be the most efficient considering architectural usage.
For AMD and Intel GPUs, only SYCL results are presented, as CUDA exclusively supports NVIDIA hardware. This limitation underscores the superior portability of SYCL compared to CUDA. The performance of the SYCL implementation on AMD and Intel GPUs presents a mixed picture. While AMD’s RX 6700 XT dGPU achieved efficiency rates comparable to NVIDIA GPUs (51.7%), the AMD Vega 6 iGPU showed considerably lower performance with only 21.3% efficiency rate. For Intel GPUs, a notable dichotomy emerged between their graphic architectures. Intel’s Gen-based iGPUs (UHD 630 and UHD 770) demonstrated impressive architectural efficiency, reaching up to 75.7%, significantly surpassing the baseline established by NVIDIA GPUs. However, Intel’s Xe-based GPUs (Arc A770 dGPU and Xe-LPG 128EU iGPU) exhibited markedly lower performance, with architectural efficiencies down to 23.4%, respectively. This suboptimal performance can be attributed to ongoing driver optimization challenges in Intel’s Arc generation, a factor that has contributed to Intel’s dGPU market share dropping below 1% in 2024 [norem24]. These performance limitations with Xe-based GPUs will likely impact the overall performance portability metrics for Intel GPUs in this study. Further investigation, including detailed code profiling, is planned to understand these performance discrepancies and their relationship to hardware architecture and driver maturity.